OpenStudy (anonymous):

A skateboarder with a mass of 45 kilograms is riding on a skateboard with a mass of 2.5 kilograms. What should be the velocity of the skateboarder if she has to attain a momentum of 264 kilogram meters/second? A. 5.5 meters/second B. 55.5 meters/second C. 1.2 × 102 meters/second D. 1.2 × 103 meters/second E. 1.2 × 104 meters/second

OpenStudy (anonymous):

5. 5 meter/sec Total mass= 45 + 2.5= 47.5 kg Total momentum = 264 momentum= mass x velocity. ie 264= 47.5 x velocity velocity = 264/47.5 = 5.5 m/sec